OpenStudy (anonymous):

a 5kg mass resting on a smooth horizontal, frictionless table is connected to a cable which passes over a pulley and then is fastened to a hanging 10kg mass. find the acceleration of the two objects and the tension of the string.

OpenStudy (anonymous):

 for m1 : m1a1 = T (1) for m2 : m2a2 = - T + m2g (2) we take (1) + (2) a(m1+m2) = m2g (a1 = a2 =a pulley no mass) we get a = 6.53m/s^2 (1) : we get T =32.67N